A Na2CO3(MM = 105.99 G/MOL) standard solution is prepared by transferring 2.4814 grams of primary standard-grade sodium carbonate to 250.0mL volumetric flask, dissolving the sample in about 100 mL of distilled deionized water and diluted to the mark. A 25.00mL aliquot is taken and titrated with 40.18mL of HCl solution. Calculate the concentration of the HCl solution.
Answer #1

molarity of Na2CO3   = W*1000/G.M.M * volume of solution

                               = 2.4814*1000/(105.99*250)   = 0.09365M

Na2CO3(aq) + 2HCl(aq) ----------------> 2NaCl(aq) + CO2(g) + H2O(l)

1 mole ----------- 2moles

Na2CO3 ------------------------------------ HCl

M1 = 0.09365M ------------------------ M2 =

V1 = 25ml --------------------------------- V2 = 40.18ml

n1 = 1 --------------------------------------- n2 = 2

            M1V1/n1   =   M2V2/n2

                M2     = M1v1n2/n1V2

                          = 0.09365*25*2/(1*40.18)   = 0.1165M >>>>answer

The concentration of HCl solution =0.1165M >>>>answer